OpenStudy (anonymous):

can someone please help me with Algerbra 2? :/

OpenStudy (anonymous):

I gotchu

OpenStudy (cggurumanjunath):

post ur q !

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

misty is very kind :) XD

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

where da question, where da at doe! XD

OpenStudy (anonymous):

Thanks everybody! :)

OpenStudy (anonymous):

what is the simplest form of the expression? 2(y-1)+8

OpenStudy (anonymous):

It's not multiple choice..

OpenStudy (anonymous):

2 * (y + 3) XD

OpenStudy (anonymous):

2y+6

OpenStudy (misty1212):

you need to "distribute the two" first \[2(y-1)=2\times y-2\times 1\]

OpenStudy (misty1212):

then add \(8\)

OpenStudy (anonymous):

good job misty but they will try to kik you out if you give the direct answer

OpenStudy (anonymous):

pplbe on here snitchin tryna get ppl kiked

OpenStudy (anonymous):

I'm not telling on anyone lol :p

OpenStudy (anonymous):

o im sorry i gave the direct answer

OpenStudy (anonymous):

no ik but ppl be on here just to do that like really bruh -_- like igreen

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

im srry @misty1212 i ment @yahman

OpenStudy (anonymous):

Awe thank you!! it feels good to get one correct!! :D

OpenStudy (anonymous):

lml wow really just pratice a lot

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

Can I ask another one?

OpenStudy (misty1212):

lol i don't think it was a direct answer the direct answer is \[2y+6\]

OpenStudy (anonymous):

yulp shoot

OpenStudy (anonymous):

another 1

OpenStudy (anonymous):

goahead

OpenStudy (anonymous):

You do the same thing to all these.. but anyways it's -b^3(3b+1)-2b^4

OpenStudy (anonymous):

simplyfy

OpenStudy (anonymous):

is that -b

OpenStudy (anonymous):

??

OpenStudy (anonymous):

−5b4−b3

OpenStudy (anonymous):

yes @Albany_Goon

OpenStudy (anonymous):

−5b4−b3

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

−5b4−b3 there

OpenStudy (anonymous):

-b^3 * (5b + 1)

OpenStudy (anonymous):

That's the answer?

OpenStudy (anonymous):

it keeps on shifting the numbers apart

OpenStudy (anonymous):

you could message it to me.

OpenStudy (anonymous):

could u 2 fan me plzzz

OpenStudy (anonymous):

ok weres the question

OpenStudy (anonymous):

y^2(3y+4)+5y^3

OpenStudy (anonymous):

@yahman

OpenStudy (anonymous):

yesh

OpenStudy (anonymous):

yes @domofan

OpenStudy (anonymous):

x^2(4x^2-3)+5x^4

OpenStudy (anonymous):

is it 18-3^2?

OpenStudy (anonymous):

3x2(3x2-1)

OpenStudy (anonymous):

thats the simplified form?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Incorrect.. that's okay though.

OpenStudy (anonymous):

wattttttttttttttt

OpenStudy (anonymous):

srryyyyyyy